What is the answers to this absolute value equation, it is asking for two answers?
2|4x-5|-2=-4
I kept trying to work out the problem but all i got was negitive numbers, and none of the answers have negitives in them.
Thanks

Answer by stanbon(75887)

2|4x-5|-2=-4
-----------
Add 2 to both sides:
---
2|4x-5| = -2
---
Divide both sides by 2:
|4x-5| = -1
----
There are NO answers because absolute
value cannot be negative.
===============================
Cheers,
Stan H.